OpenStudy (anonymous):

Given the linear equation , y=4/5x-2 find the y-coordinates of the points (-5, ), (0, ), and (5, ). Please show all of your work.

OpenStudy (anonymous):

y=(4/5)(-5)-2 y=(4/5)(0)-2 y=(4/5)(5)-2 gives y values